Equilibrium Concentrations

Consider the following reaction:
{rm{SO}}_2 {rm{Cl}}_2 left( g right); rightleftharpoons ;{rm{SO}}_2 left( g right); + ;{rm{Cl}}_2 left( g right)
K_{rm{c}} = 2.99; times ;10^{ - 7} at 227^circ rm C

If a reaction mixture initially contains 0.168 M {rm{SO}}_2 {rm{Cl}}_2, what is theequilibrium concentration of {rm{Cl}}_2 at 227 ^circ {rm C}?
Answer #1
Given that K_{rm{c}} = 2.99; times ;10^{ - 7} at227^circ rm CInitial concof {rm{SO}}_2 {rm{Cl}}_2= 0.168M{rm{SO}}_2 {rm{Cl}}_2 left( g right); rightleftharpoons ;{rm{SO}}_2 left( g right); + ;{rm{Cl}}_2 left( g right)Initial0.168M00Change-xxxEquilibrium0.168M-xxxKc = [SO2][Cl2]/[SO2Cl2]2.99 x10-7=(x)(x)/(0.168M-x)x2-2.99 x10-7(0.168M-x)=0on solving the above equation x=0.000223976M=2.23976*10-4 M∴[Cl2] atequilibrium =2.23976*10-4 M
answered by: taysha